Water at T = 30°C flows through the 400-mm-diameter concrete pipe from the reservoir at A to the one at B. Determine the flowmn=3.14
Answer #1

Given T= 300 l v. 0.801x10-6m/). d 0.4 m L = 10om e:0.8 mm = 0.8 X1903 m. TA- Zo = 3.14 +15= 18.14m. ) To find flow rate (Q)i hf 0.0234X100 Q? 1201 (0.4) ilhg=18.8930² put everything in een o ... Palm Ра Parm (7-958412 to+ ZA-ZB P9 18.8930? 2x9.81 2